# DiagnoLLM: A Hybrid Bayesian Neural Language Framework for Interpretable Disease Diagnosis

### v1

@@ Introduction
To bridge this gap, deconvolution methods have emerged that estimate CTS profiles from bulk RNA-seq using single-cell references (xu2025cell). However, existing methods often fail to generalize due to their sensitivity to reference-target mismatch, lack of robust uncertainty modeling, and inability to propagate prior biological knowledge (torroja2019digitaldlsorter). At the same time, most downstream disease classifiers treat gene expression purely as numerical input, ignoring known regulatory mechanisms such as expression quantitative trait loci (eQTLs) that offer causal, cell-type-aware insights into disease progression (nica2013expression; natri2024cell). Even when accurate classifiers are developed, their “black-box” nature severely limits clinical adoption. Physicians and patients require not only predictions, but also clear, faithful explanations grounded in known biology (blumenfeld2024cell). Recent work has explored the use of large language models (LLMs) in biomedical tasks (yang2023transformehr; gao2024raw; han2022luna; jiang2025cmfdnet), yet most applications focus on end-to-end generation or information extraction. These approaches often lack alignment with underlying model behavior, leading to hallucinations and eroding trust in high-stakes settings (omar2024utilizing; chen2025visrl; zhao2022balf; zhang2024wrong). Furthermore, LLMs exhibit well-documented limitations in handling symbolic reasoning and numerical precision (hegselmann2023tabllm; zhang2025vitcot; hu2025flowmaltrans), especially when used as primary decision-makers.
@@
We argue that a reliable clinical AI system must unify three capabilities: robust signal extraction, biologically grounded prediction, and audience-specific explanation. To this end, we propose DiagnoLLM, a hybrid neuro-symbolic framework for interpretable disease diagnosis. As shown in Figure 1, DiagnoLLM integrates Bayesian deconvolution, regulatory genomics, and LLM-based explanation into a two-stage diagnostic pipeline. Stage 1 introduces GP-unmix, a Gaussian Process-based hierarchical model that deconvolves bulk RNA-seq into CTS expression matrices using single-cell references. GP-unmix introduces posterior refinement steps to correct for reference-target shifts, incorporates uncertainty modeling via multivariate priors, and filters gene-cell-type pairs using biologically informed selection strategies. Compared to existing methods (xu2025cell; torroja2019digitaldlsorter), it achieves significantly higher gene-level recovery across tissues, species, and modalities. Stage 2 enriches prediction with eQTL-derived regulatory features that guide a two-layer neural network classifier. These regulatory priors emphasize disease-relevant transcriptional mechanisms (gusev2016integrative; natri2024cell), improving both classification performance (88.0% accuracy on AD) and biological alignment. To address the interpretability gap, we introduce a language-based reasoning module, where an LLM translates classifier outputs and feature attributions into structured, audience-specific diagnostic reports tailored for clinicians and patients. This module is grounded in attribution scores and domain knowledge, ensuring factual alignment and avoiding free-form hallucination (yu2025unified).

### v2

@@ Introduction
To bridge this gap, deconvolution methods have emerged that estimate CTS profiles from bulk RNA-seq using single-cell references (xu2025cell; tang2024cell). However, existing methods often fail to generalize due to their sensitivity to reference-target mismatch, lack of robust uncertainty modeling, and inability to propagate prior biological knowledge (torroja2019digitaldlsorter). At the same time, most downstream disease classifiers treat gene expression purely as numerical input, ignoring known regulatory mechanisms such as expression quantitative trait loci (eQTLs) that offer causal, cell-type-aware insights into disease progression (nica2013expression; natri2024cell). Even when accurate classifiers are developed, their “black-box” nature severely limits clinical adoption. Physicians and patients require not only predictions, but also clear, faithful explanations grounded in known biology (blumenfeld2024cell). Recent work has explored the use of large language models (LLMs) in biomedical tasks (yang2023transformehr; gao2024raw; han2022luna; jiang2025cmfdnet), yet most applications focus on end-to-end generation or information extraction. These approaches often lack alignment with underlying model behavior, leading to hallucinations and eroding trust in high-stakes settings (omar2024utilizing; chen2025visrl; zhao2022balf; zhang2024wrong; lin2025plan; zeng2025janusvln; zeng2025future). Furthermore, LLMs exhibit well-documented limitations in handling symbolic reasoning and numerical precision (hegselmann2023tabllm; zhang2025vitcot; hu2025flowmaltrans; lu2022understand), especially when used as primary decision-makers.
@@
We argue that a reliable clinical AI system must unify three capabilities: robust signal extraction, biologically grounded prediction, and audience-specific explanation. To this end, we propose DiagnoLLM, a hybrid neuro-symbolic framework for interpretable disease diagnosis. As shown in Figure 1, DiagnoLLM integrates Bayesian deconvolution, regulatory genomics, and LLM-based explanation into a two-stage diagnostic pipeline. Stage 1 introduces GP-unmix, a Gaussian Process-based hierarchical model that deconvolves bulk RNA-seq into CTS expression matrices using single-cell references. GP-unmix introduces posterior refinement steps to correct for reference-target shifts, incorporates uncertainty modeling via multivariate priors, and filters gene-cell-type pairs using biologically informed selection strategies. Compared to existing methods (tang2024cell; xu2025cell; torroja2019digitaldlsorter), it achieves significantly higher gene-level recovery across tissues, species, and modalities. Stage 2 enriches prediction with eQTL-derived regulatory features that guide a two-layer neural network classifier. These regulatory priors emphasize disease-relevant transcriptional mechanisms (gusev2016integrative; natri2024cell), improving both classification performance (88.0% accuracy on AD) and biological alignment. To address the interpretability gap, we introduce a language-based reasoning module, where an LLM translates classifier outputs and feature attributions into structured, audience-specific diagnostic reports tailored for clinicians and patients. This module is grounded in attribution scores and domain knowledge, ensuring factual alignment and avoiding free-form hallucination (yu2025unified).